OpenStudy (anonymous):

What is the equation (in slope-intercept form) of a line that passes through the point (3, -2) and has a slope of -2? y = 2x + 4 y = -2x + 4 y + 2 = -2(x + 3) 2x + y + 4 = 0

OpenStudy (anonymous):

first put all the questions in y=mx+b form y=2x+4 y=-2x+4 y=-2x-8 y=-2x-4

OpenStudy (anonymous):

option A is wrong because the slope needs to -2